### v1

# Condense to Conduct and Conduct to Condense

Abstract
In this paper we give the first examples of low-conductance permutations. The notion of conductance of permutations was introduced in the paper ”Indifferentiability of Confusion-Diffusion Networks” by Dodis et al., where the search for low-conductance permutations was initiated and motivated. In this paper we not only give the desired examples, but also make a general characterization of the problem – i.e. we show that low-conductance permutations are equivalent to permutations that have the information-theoretic properties of the so-called Multi-Source-Somewhere-Condensers.
1 Introduction
1.1 Background
The starting concept of this paper is the notion of permutation conductance. This notion was introduced by Yevgeniy Dodis, Martijn Stam, John Steinberger, and Tianren Liu in their seminal paper Indifferentiability of confusion-diffusion networks [11].
The above-mentioned paper concerns the theoretical analysis of the indifferentiability security of the confusion-diffusion networks, investigating how many rounds of certain constructions of block ciphers and hash functions provide adequate security. It turns out that in these considerations, the key primitive is a certain permutation, from which we expect a possibly small so-called conductance. Much more on the subject of motivation and direct relations between permutations with an appropriate conductance and the security of the studied ciphers based on confussion-diffusion networks can be found in [11]. In our work, the conductance of permutation is studied primarily as a combinatorial concept.
From the formal definition side, each permutation of the type $\pi:\{0,1\}^{nw}\to\{0,1\}^{nw}$ has a certain degree of conductance that additionally depends on a certain parameter $\alpha\in(0,1)$. This degree (denoted $\textrm{CondDeg}_{\alpha}(\pi)$) is always a real number between $1$ and $w$ and is a kind of measure of how random $\pi$ is (the more random the permutation, the closer $\textrm{CondDeg}_{\alpha}$ is to $1$; in particular, a constant permutation has $\textrm{CondDeg}_{\alpha}$ equal to $w$). It turns out that permutations with low $\textrm{CondDeg}_{\alpha}$ (”more random”) are useful in the applications described in [11]. Fortunately, the authors of [11] showed (Theorem 3, App. B, probabilistic method) that almost every random permutation has (for parameters $\alpha<\frac{1}{2}-\frac{1}{nw}$, which is sufficient for the applications described in the paper) $\textrm{CondDeg}_{\alpha}$ very close to $1$ (more precisely: it is no larger than $1+\frac{\log(3nw)}{\alpha n}$). On the other hand – there were no explicit constructions showing how such permutations could be constructed in practice.
In this paper we show for the first time direct examples of permutations with conductance degree less than $w$. Details are given in Theorem 1. Our constructions use the so-called Multi-Source-Somewhere-Condenser permutations, i.e. a certain – defined in this paper as Definition 5 – variant of classical somewhere-condensers (such as defined for exampe in Definition 1.11 in [18]).
Interestingly, we also managed to obtain the inverse result – i.e. we show that every permutation with a sufficiently low conductance degree must satisfy the Multi-Source-Somewhere-Condenser assumptions. In other words, in this paper we attempt a full characterization of the problem, which is successful for almost the entire range of parameters (see Section 3.2 for details). In particular: this means that unfortunately we cannot count on low-conductance permutations expressed as, for example, simple linear functions, as hypothesized in [11].
To conclude the introduction, it is worth noting how the considered problem shows the non-obvious interpenetration of various fields in theoretical computer science and mathematics. Namely, the original problem can be classified as an example of research in the field of theoretical analysis of symmetric ciphers (which is not common in itself, see e.g. [17, 16, 12, 8]). A nice combinatorial problem was extracted from it (and it is described as such in the original paper). Then it turned out that its analysis leaned towards information theory (in particular the theory of min-entropy extraction), so it can be seen as a work in this field (and thus among papers such as [5, 7, 20, 15, 14, 22, 19, 9, 10]). It is also worth emphasizing that other similar reasonings, i.e. solving combinatorial problems using theorems from extractor theory, are not new, this phenomenon can also be observed e.g. in [13, 21, 2, 3, 4, 1]. This all shows an insight into why this problem turned out to be surprisingly non-trivial.
1.2 Main Results
1.2.1 Main Notion
Before we present the main result of the paper, let us first quote (from [11]) the formal definition of the notion of conductance of a permutation:
Definition 1.
Consider the permutation $\pi:\{0,1\}^{nw}\to\{0,1\}^{nw}$ and the parameter $\alpha\in(0,1)$. Then $\textrm{CondDeg}_{\alpha}(\pi)$ is equal to $d$ if:
$\max_{\begin{subarray}{c}U=U_{1}\times\ldots\times U_{w}\subset(\{0,1\}^{n})^{w}\\ V=V_{1}\times\ldots\times V_{w}\subset(\{0,1\}^{n})^{w}\\ |U_{1}|=\ldots=|V_{w}|=2^{\alpha n}\end{subarray}}|\left\{(x,y)\in(U,V):y=\pi(x)\right\}|=(2^{\alpha n})^{d}.$
Since the above notion is exceptionally essential for this work, to increase its readability we add an additional, intuitive understanding of this notion below preceded by one more improtant definition:
Definition 2.
A set $X\subset(\{0,1\}^{n})^{w}$ will be called a discrete $q$-box of dimension $w$ if
$X=X_{1}\times\ldots\times X_{w},$
for some $X_{1},\ldots,X_{w}\subset\{0,1\}^{n}$ such that $|X_{1}|=\ldots=|X_{w}|=q$.
Remark: In the above definition, if the dimension is obvious from the context, we will sometimes just write that $X$ is a discrete $q$-box (omitting the phrase of dimension $w$).
Intuition of Definition 1:
In simple terms, $\textrm{CondDeg}_{\alpha}(\pi)$ determines how much a given permutation $\pi$ distorts all discrete $q$-boxes of dimension $w$ for $q=2^{\alpha n}$. More specifically, for each such $q$-box $U$ we look at its image $\pi(U)$ and look for another $q$-box $V$ such that the intersection $U\cap V$ is as large as possible. The maximum we can obtain in this way is given by $\textrm{CondDeg}_{\alpha}$ in the form of the following relation:
$\max_{U,V}(|\pi(U)\cap V|)=q^{\textrm{CondDeg}_{\alpha}}.$
In the ”worst” (least random) case we will find such $U$ that $\pi(U)$ is just a another $q$-box $V$ and then $|\pi(U)\cap V|=q^{w}$, which gives $\textrm{CondDeg}_{\alpha}=w$.
On the other hand, ”more random” permutations should intuitively behave in such a way that regardless of the choice of $U$ and $V$, the cardinality of the set $\pi(U)\cap V$ will always be much smaller than $q^{w}$. At the same time, it is easy to see that it is always possible to choose $V$ such that $|\pi(U)\cap V|$ is at least $q$. In particular, this means that we definitely have $\textrm{CondDeg}_{\alpha}(\pi)\geq 1$.
Finally, combining these intuitions and observations, we see that $\textrm{CondDeg}_{\alpha}(\pi)$ is a specific way of estimating the randomness of the permutation $\pi$, which is always a number in the range between $1$ and $w$. The more random (in the given sense) the permutation is, the smaller this factor is.
Notation.
For clarity, let us add a technical note that the original work of [11] uses a slightly different notation. Specifically, instead of the parameter $\alpha$, the parameter $q$ is used, which is simply equal to $2^{\alpha n}$ in the language of our work. Consequently, instead of $\textrm{CondDeg}_{\alpha}$, [11] refers to $\textrm{Cond}_{q}$, with the relation $\textrm{Cond}_{q}(\pi)=q^{\textrm{CondDeg}_{\alpha}(\pi)}$. Then $\textrm{Cond}_{q}(\pi)$ is a number between $q$ and $q^{w}$, not between $1$ and $w$, as is the case for $\textrm{CondDeg}_{\alpha}(\pi)$.
The above changes are dictated by a different center of gravity of our focus. The work of [11] concerns more specific applications, so the parameter $q$ (referring to the number of queries present in security analyses) is natural there. Our work clearly takes a turn towards information theory and for us the primary notion is rather the min-entropy of the source, to which $\alpha$ refers.
Results in original notation.
For consistency with the notation from the paper [11], in Appendix A we present the main results of our paper reformulated in the language of the paper [11]. Similarly, the result from [11] quoted above (Theorem 3, App. B) is originally expressed in a different language. Appendix A also contains a proof of the equivalence of both versions.
1.2.2 Main result
In this paper we give the first example of a permutation with nontrivial conductance degree:
Theorem 1.
If $n>n_{0}$ is prime and $\epsilon\leq\alpha\leq 0.8$, then there exists a permutation $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ such that
$\textrm{CondDeg}_{\alpha}(\pi)\leq w-\left\lfloor\frac{w}{3}\right\rfloor c,$
where $n_{0}=\theta\left(\frac{1}{\epsilon^{2}}\right),c=\theta(\epsilon).$
The reasoning that leads to the construction of this permutation (the final proof is in Section 2.4) requires the introduction of several lemmas (all proofs are in Section 2) and definitions, which we present below:
Definition 3.
The min-entropy of a source $X$ is defined as
$H_{\infty}(X)=-\log(\max_{x\in X}(1/P[X=x])).$
Definition 4.
A distribution $D$ over domain $X$ is a flat distribution if it is uniformly random over a set $S\subset X$. In other words: if it is uniform on its support.
We have the following well known fact on these notions:
Lemma 1.
A source $X$ with min-entropy at least $k$ is always a convex combination of flat distributions supported on sets of size $2^{k}$.
Definition 5.
The permutation $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ will be called a ($\alpha,\epsilon_{1},\epsilon_{2}$)-Multi-Source-Somewhere-Condenser if for any independent random variables $X_{1},X_{2},\ldots X_{w}\in\{0,1\}^{n}$ with min-entropy at least $\alpha n$, the distribution $\pi(X_{1},\ldots,X_{w})$ can be expressed as a convex combination of the distributions ($\sum_{1\leq i\leq w}\gamma_{i}C_{i}+\gamma R)$, such that $\gamma\leq\epsilon_{2}$ and $C_{i}]_{i}$ (i.e. the $i$-th coordinate of $C_{i}$) has min-entropy at least $(1+\epsilon_{1})\alpha n$.
Note: the definition above is based on Definition 1.11 from [18]. However we consider several sources of the input and we limit ourselves to the study of permutations. Nevertheless, its intuitive meaning remains the same: we define a transformation, which from several sources with smaller min-entropy, leads to a convex combination of distributions, each of which (except for one component of size $\gamma$) has min-entropy on some coordinate by a factor $(1+\epsilon_{1})$ larger.
Armed with the above definitions, we are ready to formulate the key technical lemma, the proof of which can be found in Section 2.1:
Lemma 2.
If the permutation $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ is a ($\alpha,\epsilon_{1},\epsilon_{2}$)-Multi-Source-Somewhere-Condenser, then $\textrm{CondDeg}_{\alpha}(\pi)$ is no larger than
$\log_{q}(q^{w-\epsilon_{1}}+\epsilon_{2}q^{w}),$
for $q=2^{\alpha n}$.
The formula in the above lemma means that if $\epsilon_{2}$ is negligable, then we will achieve $\textrm{CondDeg}_{\alpha}\approx w-\epsilon_{1}$.
So, to effectively use Lemma 2, we need a Multi-Source-Somewhere-Condenser with suitably strong parameters, which fortunately exists and may be found for example in the paper [6]. Specifically, it holds (in [6] this result and its proof are discribed as Lemma 3.14):
Lemma 3.
Consider the permutation: $\pi:\{0,1\}^{3n}\to\{0,1\}^{3n}$ (i.e. $w=3$) defined by the formula $\pi(a,b,c)\to(a,b,c+ab)$, where $a,b,c\in\{0,1\}^{n}$ are treated as elements of the field $GF(2^{n})$. Then for each $\epsilon$, there exists $\delta=\theta(\epsilon)$ such that:
If $\epsilon\leq\alpha\leq 0.8$ and $n$ is prime, then the permutation $\pi$ defined above is ($\alpha,\delta,2^{-\delta\alpha n}$)-Multi-Source-Somewhere-Condenser.
Note: the original lemma from [6] is even stronger than the formulation of Lemma 3 presented above. In fact, it is always the third coordinate (except for some set of measure $2^{-\delta\alpha n}$) that has the appropriate min-entropy, not some coordinate for different fragments of the permutation.
The corollary of Lemmas 2 and 3 is as follows:
Lemma 4.
If $\epsilon\leq\alpha\leq 0.8$, then there exists a permutation $\pi:\{0,1\}^{3n}\to\{0,1\}^{3n}$ such that
$\textrm{CondDeg}_{\alpha}(\pi)\leq 3-c\text{ (for some }c>0\text{),}$
as long as $n$ is prime and $n>n_{0}$ for some $n_{0}=\theta\left(\frac{1}{\epsilon^{2}}\right),c=\theta(\epsilon).$
The formal proof of the above lemma can be found in Section 2.3. The generalization of this lemma to higher dimensions (i.e. arbitrary $w$), which is simply the content of our main Theorem 1 – can be found in Section 2.4).
1.2.3 Converse Theorem
It turns out that the theorem also holds in the opposite direction, i.e. showing that every permutation with a suitable $\textrm{CondDeg}_{\alpha}$ is also a nontrivial Multi-Source-Somewhere-Condenser:
Theorem 2.
Let $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ be a permutation such that its conductance degree $\textrm{CondDeg}_{\alpha}(\pi)$ is less than $w(1-\epsilon_{1}-\epsilon_{2}-\epsilon_{3}-\frac{1}{\alpha n})$, for some $\epsilon_{1},\epsilon_{2},\epsilon_{3}>0$ and $0<\alpha<1$. Then $\pi$ is a ($\alpha,\epsilon_{1},r$)-Multi-Source-Somewhere-Condenser, for $r=w2^{-\alpha n\epsilon_{2}}+2^{-\alpha nw\epsilon_{3}}$.
The proof is in Section 2.2.
1.3 Intuition and Toy Examples
Consider the function $\pi_{1}(a,b,c)=(a,b,c+ab)$, where $a,b,c\in GF(2^{n})$ and $n$ is a prime number. This is of course a permutation, since for any value of $(x,y,z)$, its only inverse image is (x,y,zxy)(x,y,z-xy). We want to justify why such defined $\pi_{1}:\{0,1\}^{3n}\to\{0,1\}^{3n}$ has conductance degree $\textrm{CondDeg}_{\alpha}(\pi_{1})$ (for some fixed parameter $\alpha$) strictly less than $3$.
For this purpose, we consider two arbitrary $(2^{\alpha n})$-boxes $U=U_{1}\times U_{2}\times U_{3}$ and $V=V_{1}\times V_{2}\times V_{3}$ (i.e. $|U_{1}|=|U_{2}|=|U_{3}|=|V_{1}|=|V_{2}|=|V_{3}|=2^{\alpha n}$). Now we want (according to the Definition 1) to estimate an upper bound on $|\pi_{1}(U)\cap V|$.
This is a purely combinatorial problem, but we will look at it from the probabilistic perspective. Specifically: we will assume a uniform distribution on $U$ and then we will show some upper bound on $P_{u\leftarrow U}(\pi_{1}(u)\in V)$. This will allow us to achieve the main result, since of course we have:
$|\pi_{1}(U)\cap V|=P_{u\leftarrow U}(\pi_{1}(u)\in V)\cdot(2^{\alpha n})^{3}.$ (1)
To achieve our goal, we will use the fact that $\pi_{1}$ is a Multi-Source-Somewhere-Condenser (thus Lemma 3), and even a stronger fact (which is formulated in this way in [6]), i.e. that it is always the third (not some) coordinate of $\pi_{1}(a,b,c)$ that is the condenser, except for some negligable part of the domain. More specifically, thanks to this version of Lemma 3 we know that – since we are considering a uniform distribution on $U$ – and thus we have that $U_{1},U_{2}$ and $U_{3}$ are independent and have min-entropy $\alpha n$, there exists such $\delta=\theta(\alpha)$ that $\pi_{1}(U)$ can be written as a convex combination of two distributions $C_{3}$ and $R$ (that is, $\pi(U)=\gamma_{3}C_{3}+\gamma R$ for some $\gamma_{3},\gamma>0,\gamma_{3}+\gamma=1$) with the following properties:
$\gamma\leq 2^{-\delta\alpha n}$ (2)
and $C_{3}]_{3}$ (i.e. the third coordinate of $C_{3}$) has min-entropy at least $(1+\delta)\alpha n$. This last fact can also be formally written as follows: for any $x\in\{0,1\}^{n}$, it holds
$P_{v\leftarrow C_{3}}(v\in\{0,1\}^{n}\times\{0,1\}^{n}\times\{x\})\leq 2^{-(1+\delta)\alpha n}.$ (3)
Let us additionally denote $V_{3}=\{x_{1},\ldots,x_{2^{\alpha n}}\}$ and we can finally complete the main estimation:
[…]
Hence:
[…]
This of course means that $\textrm{CondDeg}_{\alpha}\leq 3-\delta+\frac{1}{\alpha n}$, which is strictly smaller than $3$ for sufficiently large $n$, so $\pi_{1}$ really has nontrivial conductance.
A completely analogous reasoning can be carried out for the permutation $\pi_{2}(a,b,c)=(a,b+ac,c)$, establishing that and $\pi_{2}$ is also a permutation with nontrivial conductance.
Going even further, let us now consider the following permutation:
$\pi_{3}(a,b,c)=\begin{cases}(a,b,c+ab)&\text{for }a\text{ even},\\ (a,b+ac,c)&\text{for }a\text{ odd},\end{cases}$
which, as we can see, is a kind of mix of the permutations $\pi_{1}$ and $\pi_{2}$, so it is certainly – as can be checked from the definition – a Multi-Source-Somewhere-Condenser. Reasoning that is only slightly more complicated than the one given above (and thus similar to the proof of Lemma 2) allows us to prove that $\pi_{3}$ is also a permutation with nontrivial conductance.
Interestingly and importantly, in this work we managed to show that specifying more complex constructions is not possible at all. That is: every permutation with nontrivial conductance must have a character exactly like $\pi_{3}$ – so it must be a ”mix” of different concrete condensers on different coordinates (i.e. it must be a Somewhere-Condenser). This fact is described in detail in Theorem 2.
2 Technical Part: formal proofs
2.1 Proof of Lemma 2
(This proof is essentially a direct generalization of the reasoning presented in Section 1.3.)
To prove the statement, we need to consider two arbitrary $(2^{\alpha n})$-boxes $U=U_{1}\times\ldots\times U_{w},V=V_{1}\times\ldots\times V_{w}$ (i.e. $|U_{1}|=\ldots=|U_{w}|=|V_{1}|=\ldots=|V_{w}|=2^{\alpha n}$) and show that $|\pi(U)\cap V|\leq(2^{\alpha n})^{w-\epsilon_{1}}+\epsilon_{2}(2^{\alpha n})^{w}$. Equivalently, we will consider the uniform distribution on $U$ and prove that
$P_{u\leftarrow U}(\pi(u)\in V)\leq 2^{-\epsilon_{1}\alpha n}+\epsilon_{2}.$
Since $U_{1},\ldots,U_{w}$ are independent and have min-entropy $\alpha n$, we can use the fact that $\pi$ is a $(\alpha,\epsilon_{1},\epsilon_{2})$-Multi-Source-Somewhere-Condenser. We then get that
$\pi(U)=\sum_{i=1}^{w}\gamma_{i}C_{i}+\gamma R,$
where $\gamma\leq\epsilon_{2}$ and $C_{i}]_{i}$ (the $i$-th coordinate of $C_{i}$) has min-entropy of at least $(1+\epsilon_{1})\alpha n$.
This last fact can be more conveniently rewritten as follows: for any $x\in\{0,1\}^{n}$, it holds
$P_{v\leftarrow C_{i}}(v\in\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{x\}}\times\ldots\times\{0,1\}^{n})\leq 2^{-(1+\epsilon_{1})\alpha n}.$
If we additionally denote $V_{i}=\{x_{1},\ldots,x_{2^{\alpha n}}\}$, then from the above we also have:
$\begin{split}P_{v\leftarrow C_{i}}(v\in V)&\leq\sum_{j=1}^{2^{\alpha n}}P_{v\leftarrow C_{i}}(v\in\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{x_{j}\}}\times\ldots\times\{0,1\}^{n})\leq\\ &\leq 2^{\alpha n}\cdot 2^{-(1+\epsilon_{1})\alpha n}=2^{-\epsilon_{1}\alpha n}.\end{split}$ (4)
Given the above, we can proceed to the main estimation:
[…]
which completes the proof.
2.2 Proof of Theorem 2
We are to show that $\pi$ is a $(\alpha,\epsilon_{1},r)$-Multi-Source-Somewhere-Condenser. By definition, we should consider some independent random variables $U_{1},\ldots,U_{w}$ such that their min-entropy is at least $\alpha n$ and show that the distribution $\pi(U_{1},\ldots,U_{w})$ has the appropriate properties. However, without loss of generality (using Lemma 1), we can assume that the distribution of $U=U_{1}\times\ldots\times U_{w}$ is flat, that is, it is really just a uniformly distributed discrete $(2^{\alpha n})$-box $U$ (as defined in Definition 2). Then our original statement (after translating to the discrete interpretation) translates into the following statement:
Claim 1.
The set $\pi(U)$ can be partitioned into disjoint sets $\{C_{i}\}_{i=1}^{w}$ and $R$ such that:
  • $|R|\leq r\cdot 2^{\alpha nw}$, where $r=w2^{-\alpha n\epsilon_{2}}+2^{-\alpha nw\epsilon_{3}}$;
  • for each $C_{i}$ and for each $y\in\{0,1\}^{n}$ we have:
    $|C_{i}\cap\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{y\}}\times\ldots\times\{0,1\}^{n}|\leq 2^{-(1+\epsilon_{1})\alpha n}\cdot|C_{i}|.$
Notational note: in the above statement $R$ and $C_{i}$ simply denote subsets of $\{0,1\}^{wn}$, while in the original Definition 5 the same letters would denote uniform distributions on these sets.
Besides, let us remind that we know that the set $\pi(U)$ has cardinality $2^{\alpha nw}$ and (thanks to the assumption that $\pi$ has the appropriate degree of conductance) that for any other $(2^{\alpha n})$-box $V$ the following always holds:
$|\pi(U)\cap V|<2^{\alpha nw(1-\epsilon_{1}-\epsilon_{2}-\epsilon_{3}-\frac{1}{\alpha n})}.$ (5)
Before we proceed to the direct proof of Claim 1 (that is, giving the appropriate way of partitioning $\pi(U)$), we need one more auxiliary definition:
Definition 6.
We say that the set $X$ has at point $y\in\{0,1\}^{n}$, an $i$-Battleneck-Slice
$B=X\cap\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{y\}}\times\ldots\times\{0,1\}^{n}$
, if it holds
$0<|B|<2^{\alpha n(w-1-\epsilon_{1}-\epsilon_{2})}.$
Intuitively: $i$-Battleneck-Slice at point $y$ means that the projection of $X$ onto the $i$-coordinate at point $y$ is nonempty, but suitably thin.
Now the procedure for partitioning the set $\pi(U)$ is as follows.
Procedure for partitioning $\pi(U)$:
  1. Step 1:
    First, we initialize the auxiliary variables: for each $i$, let $\tilde{C_{i}}$ be the empty set and let $\tilde{R}:=\pi(U)$.
  2. Step 2:
    If $\tilde{R}$ contains any (for any $i$ and at any point $y$) $i$-Battleneck-Slice $B$, perform the assignments:
    $\begin{split}\tilde{R}&:=\tilde{R}\setminus B;\\ \tilde{C_{i}}&:=\tilde{C_{i}}\cup B.\end{split}$
  3. Step 3:
    Loop Step 2 until $\tilde{R}$ no longer contains any Battleneck-Slice.
  4. Step 4:
    Assign $R_{0}:=\tilde{R}$ and $R_{1}:=\emptyset$.
  5. Step 5:
    For each $i$, if $|\tilde{C_{i}}|>2^{\alpha n(w-\epsilon_{2})}$, then assign $C_{i}:=\tilde{C_{i}}$, otherwise: assign $C_{i}:=\emptyset$ and $R_{1}:=R_{1}\cup\tilde{C_{i}}$.
  6. Step 6:
    Assign $R:=R_{0}\cup R_{1}$.
Intuitively: from the set $\pi(U)$ we cut off all Battleneck-Slices one by one until we are left with $R_{0}$, which no longer contains sets of this type. Then all cut off $i$-Battleneck-Slices concerning the $i$-th coordinate together form $C_{i}$, unless their measure is sufficiently small, in which case $C_{i}$ is empty and these Battleneck-Slices go to $R_{1}$. The final $R$ is the sum of $R_{0}$ and $R_{1}$.
It remains to analyze that the partitioning constructed in this way actually satisfies the requirements of Claim 1. So we have both necessary analyses in turn:
Analysis of $C_{i}$.
If $C_{i}$ is empty, then the statement trivially holds. Otherwise, it must be
$|C_{i}|>2^{\alpha n(w-\epsilon_{2})}.$
Additionally, for any $y$, it holds
$|C_{i}\cap\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{y\}}\times\ldots\times\{0,1\}^{n}|<2^{\alpha n(w-1-\epsilon_{1}-\epsilon_{2})},$
since every such set as above on the left is either empty or (as follows from the construction) equal to exactly one $i$-Battleneck-Slice cut off from $\tilde{R}$.
Combining these facts, we can conclude that:
$\begin{split}|C_{i}\cap\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{y\}}\times\ldots\times\{0,1\}^{n}|&<2^{\alpha n(w-1-\epsilon_{1}-\epsilon_{2})}=\\ &={2^{\alpha n(-1-\epsilon_{1})}}\cdot{2^{\alpha n(w-\epsilon_{2})}}<\\ &<{2^{-(1+\epsilon_{1})\alpha n}}\cdot|C_{i}|,\end{split}$
what was needed to be shown.
Analysis of $R$.
We have to show $|R|\leq r\cdot 2^{\alpha nw}\text{, where }r=w2^{-\alpha n\epsilon_{2}}+2^{-\alpha nw\epsilon_{3}},$ that is,
$|R|\leq w2^{(w-\epsilon_{2})\alpha n}+2^{(1-\epsilon_{3})\alpha nw}.$
Additionally, we know that $R=R_{0}\cup R_{1}$. So it is enough to show two inequalities:
$\begin{split}|R_{1}|&\leq w\cdot 2^{(w-\epsilon_{2})\alpha n},\\ |R_{0}|&\leq 2^{(1-\epsilon_{3})\alpha nw}.\end{split}$
The first one is a trivial consequence of the partitioning procedure: $R_{1}$ in the worst case could have had $w$ times added the set $\tilde{C_{i}}$ of cardinality at most $2^{\alpha n(w-\epsilon_{2})}$, which is exactly the postulated bound.
The last inequality follows directly from the fact that $R_{0}$ does not contain any Battleneck-Slice, the bound $|R_{0}|\leq|\pi(U)|=2^{\alpha nw}$ and the property (5) (this property holds for $\pi(U)$, so obviously also for $R_{0}\subset\pi(U)$), which, however, requires a slightly deeper analysis, which (thus concluding the entire proof) we present below:
Claim 2.
If the set $R_{0}$ does not contain any Battleneck-Slice, $|R_{0}|\leq 2^{\alpha nw}$ and
$|R_{0}\cap V|<2^{\alpha nw(1-\epsilon_{1}-\epsilon_{2}-\epsilon_{3}-\frac{1}{\alpha n})}$ (6)
for any discrete $2^{\alpha n}$-box $V$ of dimension $w$, then
$|R_{0}|\leq 2^{(1-\epsilon_{3})\alpha nw}.$
Reminder: the formal definition of a discrete box is given in Definition 2.
To prove the above statement, let us assume for the sake of contradiction that $|R_{0}|>2^{(1-\epsilon_{3})\alpha nw}$. Then:
First, let us note that $R_{0}$ is contained in a discrete $(2^{(1+\epsilon_{1}+\epsilon_{2})\alpha n})$-box $V_{big}$. This is actually the case because $|R_{0}|\leq 2^{\alpha nw}$ and $R_{0}$ does not contain any Battleneck-Slice, so at each coordinate it takes at most:
$\frac{2^{\alpha nw}}{2^{(w-1-\epsilon_{1}-\epsilon_{2})\alpha n}}=2^{(1+\epsilon_{1}+\epsilon_{2})\alpha n}$
different values.
This means that $V_{big}$ is a sum of at most
$\left(\left\lceil\frac{2^{(1+\epsilon_{1}+\epsilon_{2})\alpha n}}{2^{\alpha n}}\right\rceil\right)^{w}=(\lceil 2^{(\epsilon_{1}+\epsilon_{2})\alpha n}\rceil)^{w}\leq(2^{(\epsilon_{1}+\epsilon_{2})\alpha n}+1)^{w}\leq(2^{(\epsilon_{1}+\epsilon_{2})\alpha n+1})^{w}$
discrete $(2^{\alpha n}$)-boxes.
Since $|R_{0}|>2^{(1-\epsilon_{3})\alpha nw}$, at least one of these boxes contains
$\frac{2^{(1-\epsilon_{3})\alpha nw}}{(2^{(\epsilon_{1}+\epsilon_{2})\alpha n+1})^{w}}=2^{\alpha nw(1-\epsilon_{3}-\epsilon_{1}-\epsilon_{2}-\frac{1}{\alpha n})}$
points, which, however, contradicts the (6) property, and thus completes the proof.
2.3 Proof of Lemma 4
We simply apply Lemma 2 to the permutation from Lemma 3:
That is, first from Lemma 3 we have that there exists a permutation $\pi:\{0,1\}^{3n}\to\{0,1\}^{3n}$, which is a ($\epsilon,\delta,2^{-\delta\alpha n}$)-Multi-Source-Somewhere-Condenser for some $\delta=\theta(\epsilon)$. Now, additionally denoting $q=2^{\alpha n}$, from Lemma 2 we have that this permutation has $\textrm{CondDeg}_{\alpha}(\pi)$ such that:
$\begin{split}\textrm{CondDeg}_{\alpha}(\pi)&\leq\log_{q}(q^{3-\delta}+2^{-\alpha\delta n}\cdot q^{3})=\\ &=\log_{q}(q^{3-\delta}+q^{-\delta}\cdot q^{3})=\log_{q}(2\cdot q^{3-\delta})=\\ &=\log_{q}{q^{3-\delta+\frac{1}{\log q}}}=3-\delta+\frac{1}{\log q}=\\ &=3-c,\end{split}$
for $c=\delta-\frac{1}{\log q}=\delta-\frac{1}{\alpha n}$.
Finally, to achieve the postulated result $c=\theta(\alpha)=\theta(\delta)$, it is enough to choose the parameters so that:
$\delta\geq c\geq\delta/2.$
The left inequality follows trivially from the formula on $c$ obtained above. To obtain the right one, we want
$\frac{1}{\alpha n}\leq\delta/2,$
which means
$n\geq\frac{2}{\delta\alpha}=O\left(\frac{1}{\epsilon^{2}}\right),$
and is true by assumption so completes the proof.
2.4 Final Proof of Theorem 1
The final proof is a direct consequence of Lemma 4, since our final construction for dimension $w$ will be just a simple serial repetition of many copies of the dimension $3$ permutation $\pi$ from the aforementioned lemma.
To formalize this construction, let us first denote the permutation $\pi:\{0,1\}^{3n}\to\{0,1\}^{3n}$ (coming from Lemma 4) on successive coordinates as follows:
$\pi(x,y,z)=(\pi_{1}(x,y,z),\pi_{2}(x,y,z),\pi_{3}(x,y,z))\text{, where }x,y,z\in\{0,1\}^{n}.$
Now let $w\geq 3$ and the permutation $\pi^{(w)}:\{0,1\}^{wn}\times\{0,1\}^{wn}$ be defined as follows (for now let’s assume that $w$ is a multiple of $3$):
$\begin{split}\pi^{(w)}(x_{1},\ldots,x_{w})=\Big{(}&\pi_{1}(x_{1},x_{2},x_{3}),\pi_{2}(x_{1},x_{2},x_{3}),\pi_{3}(x_{1},x_{2},x_{3}),\\ &\pi_{1}(x_{4},x_{5},x_{6}),\pi_{2}(x_{4},x_{5},x_{6}),\pi_{3}(x_{4},x_{5},x_{6}),\\ &\ldots\\ &\pi_{1}(x_{w-2},x_{w-1},x_{w}),\pi_{2}(x_{w-2},x_{w-1},x_{w}),\pi_{3}(x_{w-2},x_{w-1},x_{w})\Big{)}.\end{split}$
If $w$ leaves a remainder $1$ when divided by $3$, we define:
$\pi^{(3k+1)}(x_{1},\ldots,x_{3k+1})=(\pi^{(3k)}(x_{1}\ldots,x_{3k}),x_{3k+1}),$
and similarly, when $w$ divided by $3$ leaves a remainder $2$:
$\pi^{(3k+2)}(x_{1},\ldots,x_{3k+2})=(\pi^{(3k)}(x_{1}\ldots,x_{3k}),x_{3k+1},x_{3k+2}).$
For example:
$\begin{split}\pi^{(7)}(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6},x_{7})=\Big{(}&\pi_{1}(x_{1},x_{2},x_{3}),\pi_{2}(x_{1},x_{2},x_{3}),\pi_{3}(x_{1},x_{2},x_{3}),\\ &\pi_{1}(x_{4},x_{5},x_{6}),\pi_{2}(x_{4},x_{5},x_{6}),\pi_{3}(x_{4},x_{5},x_{6}),x_{7}\Big{)}.\end{split}$
It remains to estimate $\textrm{CondDeg}_{\alpha}(\pi^{(w)})$ of the presented construction. For this purpose, we consider two arbitrary $(2^{\alpha n})$-boxes: $U=U_{1}\times\ldots\times U_{w}$ and $V=V_{1}\times\ldots\times V_{w}$ (i.e. $|U_{1}|=\ldots|U_{w}|=|V_{1}|=\ldots|V_{w}|=2^{\alpha n}$) and we estimate $|\pi^{(w)}(U)\cap V|=(2^{\alpha n})^{\textrm{CondDeg}_{\alpha}(\pi^{(w)})}$.
As in Sections 1.3 and 2.1, it will be more convenient to assume a uniform distribution on $U$ and use the probabilistic interpretation:
$(2^{\alpha n})^{\textrm{CondDeg}_{\alpha}(\pi^{(w)})}=|\pi^{(w)}(U)\cap V|=P_{u\leftarrow U}(\pi^{(w)}(u)\in V)\cdot 2^{\alpha nw}.$
Before we move on to the main estimation, let us just recall that for any $X,Y,Z,A,B,C\subset\{0,1\}^{n}$ such that $|X|=|Y|=|Z|=|A|=|B|=|C|=2^{\alpha n}$ we have (from Lemma 4 and again the probabilistic interpretation of the inequality $\textrm{CondDeg}_{\alpha}(\pi)\leq 3-c$):
$P_{(x,y,z)\leftarrow X\times Y\times Z}(\pi(x,y,z)\in(A,B,C))\cdot(2^{\alpha n})^{3}\leq(2^{\alpha n})^{3-c}.$ (7)
Armed with this fact, we can finally move on to the main estimation:
[…]
where $(\star)$ follows immediately from the structure of the construction of $\pi^{(w)}$ (primarily from the fact that successive triples in the construction are fully independent of each other).
Finally, from (8) we have that $\textrm{CondDeg}_{\alpha}(\pi^{(w)})\leq w-\left\lfloor\frac{w}{3}\right\rfloor c$, which completes the proof.
3 Open Problems and Future Work
In this section we list a number of issues that remain open and may be worth further investigation.
3.1 Wider Range of Parameters
The main result (Theorem 1) largely relies on Lemma 3.14 of [6] (i.e. Lemma 3 in this paper). It contains restrictions that – we suspect – can be avoided. First of all, we ask: whether the same result can be obtained for $n$, which are not prime numbers, and also whether the parameters $n$ and $c$ must really depend on $\epsilon$ (the work of [6] even gives hope for removing these dependencies – one would like to simply use Lemma 3.1 from that paper, but unfortunately in our solution it turned out that we were able only to use Lemma 3.14, which is a version of Lemma 3.1 with weaker assumptions and a weaker thesis.).
Additionally, there remains an open question about low-conductance permutations for $w=2$.
3.2 The Converse Theorem for Weak Conductance
The Converse Theorem (Theorem 2) applies to almost all cases, except when
$w\left(1-\frac{1}{n}\right)<\textrm{CondDeg}_{\alpha}<w.$
It remains to be investigated.
3.3 Always-Condenser should be better then Somewhere-Condenser
Recall that our solution for the dimension $3$ is the following:
$\pi_{1}(a,b,c)=(a,b,c+ab).$
While we can easily believe that the third coordinate of $\pi_{1}$ really behaves very randomly, since it is a condenser, it seems intuitive that the following construction should have even lower conductance then $\pi_{1}$ (i.e. both the second and third coordinates are condensers):
$\pi_{2}(a,b,c)=(a,ba+c,ca-b).$
However, we cannot prove formally that $\pi_{2}$ indeed is better.
(For clarity: we can show that when $n$ is a prime such that $n\equiv 3\mod 4$, then $\pi_{2}$ is indeed a permutation; in a more complicated way, we can improve this example by constructing a permutation where all three coordinates will simultaneously have the form of a condenser.)
The above question can be naturally generalized for higher dimensions. In other words: we hypothesize that if all coordinates (and not, for example, every third one like in our construction) of a permutation are condensers, then such a permutation has an even better conductance degree than the construction that derives from Lemma 2.
Appendix A Original language formulations from [11]
A.1 Theorem 3, Appendix B from [11]
The result quoted in the introduction:
$\textrm{CondDeg}_{\alpha}\leq 1+\frac{\log(3nw)}{\alpha n}\text{, provided }\alpha\leq\frac{1}{2}-\frac{1}{nw},$
was originally formulated as follows:
$\textrm{Cond}_{q}\leq 2nqw+1\text{, provided }q^{2w}\leq 2^{nw}/4.$
It is of course a consequence of the original result, because on the one hand we have (equivalence of assumptions):
$q^{2w}\leq 2^{nw}/4\iff 2^{2\alpha nw}\leq 2^{nw-2}\iff 2\alpha nw\leq nw-2\iff\alpha\leq\frac{1}{2}-\frac{1}{nw},$
and on the other hand we have:
$\begin{split}\textrm{CondDeg}_{\alpha}&=\log_{q}(\textrm{Cond}_{q})\leq log_{q}(2nqw+1)\leq log_{q}(3nqw)=\\ &=1+\log_{q}(3nw)=1+\frac{\log(3nw)}{\log(q)}=1+\frac{\log(3nw)}{\alpha n}.\end{split}$
A.2 Main results of the paper in the notation from the paper [11]
For the sake of clarity, we present the main results once again, this time rewritten in the original notation from the paper [11]:
Theorem 1’.
If $n>n_{0}$ is a prime number and $\epsilon\leq\frac{\log q}{n}\leq 0.8$, then there exists a permutation $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ such that:
$\textrm{Cond}_{q}(\pi)\leq q^{w-\left\lfloor\frac{w}{3}\right\rfloor c},$
where
$n_{0}=\theta\left(\frac{1}{\epsilon^{2}}\right),c=\theta(\epsilon).$
Theorem 2’.
Let $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ be a permutation such that $\textrm{Cond}_{q}(\pi)<q^{w(1-\epsilon_{1}-\epsilon_{2}-\epsilon_{3}-\frac{1}{\log q})}$, for some $\epsilon_{1},\epsilon_{2},\epsilon_{3}>0$. Then $\pi$ is a ($\frac{\log q}{n},\epsilon_{1},wq^{-\epsilon_{2}}+q^{-w\epsilon_{3}}$)-Multi-Source-Somewhere-Condenser.